OpenStudy (aaronandyson):

The point P is the foot of perpendicular from A(-5,7) to the line whose equation is 2x - 3y =+ 18 =0 Determine: 1.the equation of line AP 2.the coordinates of P @jayzdd

OpenStudy (anonymous):

we can first solve for y 2x - 3y = 18 2x - 18 = 3y y = 2/3*x - 6

OpenStudy (aaronandyson):

Slope of y = 2/3

OpenStudy (aaronandyson):

what next?

OpenStudy (anonymous):

so we know the given line, and teh point A(-5,7) this is what we have so far graphically https://www.desmos.com/calculator/x2ttl2ydj7

OpenStudy (anonymous):

1. equation of line AP we know the line is perpindicular, so the slope will be -3/2 and it goes through point (-5,7) y - 7 = -3/2 * (x - (-5) ) simplify this

OpenStudy (anonymous):

1. equation of line AP we know the line is perpindicular, so the slope will be -3/2 and it goes through point (-5,7) y - 7 = -3/2 * (x - (-5) ) simplify this here is picture of perpindicular line https://www.desmos.com/calculator/z5s2sepstl

OpenStudy (anonymous):

y = -3/2*x - 15/2 + 7

OpenStudy (anonymous):

y = -3/2*x - 15/2 + 7 y = -3/2*x - 1/2 Now find where the two lines intersect, that will be your foot P

OpenStudy (anonymous):

y = -3/2*x - 15/2 + 7 y = -3/2*x - 1/2 Now find where the two lines intersect, that will be your foot P solve \( y_1 = y_2\) 2/3*x - 6 = -3/2*x - 1/2

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

2/3*x + 3/2*x = 6 - 1/2 13/6 x = 11/2 x = 11/2 * 6 /13 = 33/13 the x value is 33/13 y value is -56/13